Suppose that the length of a rectangle with an area of 24 square units is unknown. Let x
represent the length. Write an expression for the width in terms of x. Use this expression to
write a formula for the perimeter, p, in terms of the unknown value, x.
1
Expert's answer
2012-09-25T11:03:45-0400
Let's denote the area of a rectangle by A and its width by w. Then

A = w*x.

So, the expression for the width is

w = A/x.

The formula for the perimeter is

p = 2(x+w) = 2(x+A/x).